the doctor orders a 14 1/2 % solution he asks the pharmacy to prepare 35 ml of the solution how many grams of pure drug will be in 35 ml of this 14 1/2 % solution round to the nearest tenth
Answer by stanbon(75887)

the doctor orders a 14 1/2 % solution
he asks the pharmacy to prepare 35 ml of the solution
how many grams of pure drug will be in 35 ml of this 14 1/2 % solution round to the nearest tenth.
------------
Question:: Are you mixing the 35% solution with pure water?
-----------
Assuming you are you get:
active + active = active
1*x + 0(35-x) = 0.145*35
------
Multiply thru by 1000 to get:
1000x + 0 = 145*35
x = 5.075 ml (amount of pure drug needed)
35-x = 29.925 ml (amout of pure water needed)
-------------------
Cheers,
Stan H.
----------------